40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 2 15:14:54 2012, jenne, jamie, Update, IOO, PZT diagnosis 
    Reply  Tue Oct 2 16:32:43 2012, Jenne, Update, IOO, PZT diagnosis 
       Reply  Thu Oct 4 01:06:52 2012, Jenne, Update, IOO, PZT diagnosis 
          Reply  Mon Oct 8 14:19:17 2012, Jenne, Update, IOO, PZT diagnosis - not fixed yet 
             Reply  Mon Oct 8 18:45:48 2012, Jenne, Update, IOO, PZT diagnosis - not fixed yet, possible solution 
                Reply  Mon Oct 8 21:42:17 2012, Jenne, Update, IOO, PZT diagnosis - not fixed yet, possible solution 
                Reply  Tue Oct 9 00:25:33 2012, Jenne, Update, IOO, PZTs - hacky solution in place!! 
Message ID: 7509     Entry time: Tue Oct 9 00:25:33 2012     In reply to: 7505
Author: Jenne 
Type: Update 
Category: IOO 
Subject: PZTs - hacky solution in place!! 

[Evan, Jenne]

We applied some volts across both the pitch and yaw pin sets of the ribbon cable that goes to PZT2.  We ended up with ~40V yaw and ~14.5V pitch.  That was the nice happy center of the clipping that we can see on the AS camera.  Once we found the center of the PZT clipping range with the ITMY beam, we recentered the AS camera (actually, this took a few iterations, but now it's good). 

We then aligned MICH, but aren't able to get it to lock.  Before falling asleep, we have decided to align the PRM and SRM, so right now we have a flashing DRMI.  Both the SRMI and PRMI look a little funny the closer you get to 'good' alignment, so I'll investigate a little more tomorrow, and include pictures.  (The video capture script has barfed again, and I'm not in the mood to deal with it today.)

ELOG V3.1.3-